# Representing in Low Rank I: conjugacy, topological and homological aspects

### v1

@@ 7.2. Somme illustrative examples
If $\Gamma$ is nilpotent, then it even satisfies the 33rd Zassenhaus conjecture [WeissPGroup, WeissNilpotent]. The only non-nilpotent $\Gamma\in\mathcal{S}(\operatorname{M}_{2}({\mathbb{Z}}[\sqrt{-1}]))$ with $|\Gamma|\neq 48$ or $96$, is the metacyclic group $C_{3}\rtimes C_{8}$ (as can be seen from Table˜B in Appendix˜A). Now note that all strict subgroups of $C_{3}\rtimes C_{8}$ are cyclic. As $C_{3}\rtimes C_{8}$ is metacyclic it satisfies the 11st Zassenhaus conjecture [HertCC, CMdR]. Moreover by Whitcomb’s theorem it also satisfies the isomorphism problem, thus altogether indeed has the subgroup isomorphism property for all its subgroups, as claimed.
@@
Now the only finite subgroups of $\operatorname{GL}_{2}({\mathbb{Z}}[\sqrt{-1}])$ which are not isomorphic to a subgroup of $\mathrm{U}_{2}({\mathbb{F}}_{3})$ are $D_{12}$ and $Dic_{3}$. Thus by part (1), it remains to show that $He\ncong D_{12}$ or $Dic_{3}$. Now if $He\cong D_{12}$, then this would imply that $V({\mathbb{Z}}[Ge])$, with $Ge\cong\mathrm{U}_{2}({\mathbb{F}}_{3})$, contains a subgroup isomorphic to $D_{6}$. However, the latter is a Frobenius group and thus by [PdRV, Theorem D], $Ge$ would also contain a copy of $D_{6}$, a contradiction. The dicyclic case is treated similarly. ∎
@@ 7.3.1. The setting of components of group rings
By (7.7) we have that $\Phi_{e}(H)$ is a finite subgroup of $V(R[Ge])$, and therefore $|\Phi_{e}(H)|\mid|Ge|$, see for example [AngelSurvey, Corollary 2.7]. To prove (i), we show that $|He|\mid|\Phi_{e}(H)|$. For this, note that (7.8) implies that $He=\pi_{e}(H)=\sigma_{e}(\Phi_{e}(H))$ is an epimorphic image of $\Phi_{e}(H)$ and hence $|He|$ divides $|\Phi_{e}(H)|$, as desired. Since $V(R[Ge])$ has exponent $\exp(Ge)$ by (7.3), it follows moreover that $\exp(\Phi_{e}(H))\mid\exp(Ge)$ (since $\Phi_{e}(H)\leqslant V(R[Ge])$ and the fact that the exponent of a subgroup divides the exponent of the overlying group). But since $He$ is an epimorphic image of $\Phi_{e}(H)$, it follows that $\exp(He)\mid\exp(\Phi_{e}(H))$, which finishes (ii).
@@
Let $G$ and $H$ be a finite groups such that ${\mathbb{Z}}G\cong{\mathbb{Z}}H$ as rings. Then $Ge\cong He$ for every $e\in\operatorname{PCI}({\mathbb{Q}}G)$.
@@ 7.3.3. The setting of simple algebras of reduced degree 2
Let $D$ be an arbitrary finite-dimensional division $\mathbb{Q}$}-algebra. It will be useful to distinguish primitive and imprimitive groups. Recall the following definition (see [Banieqbal, section 2.1]).
@@
First we consider the case that $\mathcal{Z}(D)={\mathbb{Q}}$. Then $B$ is a simple $\mathcal{Z}(D)$-subalgebra of $\operatorname{M}_{2}(D)$. Therefore, the double centraliser theorem [EricAngel1, Theorem 2.10.1] implies that $\dim_{{\mathbb{Q}}}(B)$ divides $\dim_{{\mathbb{Q}}}\operatorname{M}_{2}(D)$. The latter in turn is a divisor of $16$ since $\operatorname{M}_{2}(D)$ is exceptional. Now recall that $\dim_{\mathcal{Z}(B)}(B)$ is a square and the reduced degree of $B$ is less than or equal to $2$ by [Banieqbal, pg 455, Lemma]. Together with the assumptions on $B$, and the fact that
@@ 7.4. Finite subgroups of exceptional simple algebras
Amitsur [Amitsur] classified all finite subgroups of a finite-dimensional division algebra. Moreover, for each of these groups, he described an explicit division algebra containing it. From this, one can in particular deduce the possible finite spanning subgroups of a given division algebra. In a highly technical work, Banieqbal [Banieqbal] described the primitive spanning subgroups of $\operatorname{M}_{2}(D)$. However, in case of an exceptional (simple) matrix algebra, a direct and short proof was provided in [EKVG] for the finite groups spanning such simple algebras. In this section we also describe the non-spanning finite subgroups.
@@ 7.4.1. The imprimitive subgroups
We start by describing the imprimitive groups. For the following statement, we follow the techniques from the proof of [Banieqbal, Lemma 2.2].
@@
Hence in particular the finite subgroups of $\operatorname{GL}_{2}(D)$ of imprimitive type are easy to describe in terms of finite subgroups of ${\mathcal{U}}(D)$. We will mainly be interested in finite subgroups of $\operatorname{GL}_{2}(\mathcal{O})$. By Lemma˜7.21 and [EKVG], if $\operatorname{M}_{2}(D)$ is exceptional, we have that
@@
Moreover, the above division algebras are left norm euclidean, see [BJJKT, Remarks 3.13 & 6.14], and hence have up to conjugation a unique maximal order [CeChLez, Proposition 2.9]. The unique maximal order of $\mathbb{Q}$} is $\mathbb{Z}$}, which has unit group $C_{2}=\{\pm 1\}$. For ${\mathbb{Q}}(\sqrt{-d})$ it is its ring of integers, denoted $\mathcal{I}_{d}$ whose unit group is as following, see [BJJKT, Section 3]:
@@
This follows from a straightforward case-by-case analysis for each isomorphism type of exceptional components $A$ such that $\operatorname{GL}_{2}(D)={\mathcal{U}}(A)$, combined with the list of spanning subgroups given by [EricAngel1, Corollary 12.12, Tab. 12.1 and Tab. 12.2] (see also Table˜B in Appendix˜A). By Lemma˜7.28 (see also remark˜7.29), all finite imprimitive subgroups are isomorphically contained in a largest imprimitive subgroup, which is of the form $({\mathcal{U}}(\mathcal{O})\times{\mathcal{U}}(\mathcal{O}))\rtimes C_{2}$. We summarise the maximal imprimitive subgroups which appear as a subgroup of a spanning group in Table˜1 by giving the SmallGroupID of a spanning group containing it. Lemma˜7.28 then allows to deduce all other imprimitive subgroups which are contained in a spanning group. Note that from Table˜1, it follows that only for $\operatorname{M}_{2}(\mathbb{H}_{5})$, there are no order-admissible spanning groups for the largest imprimitive subgroup $({\mathcal{U}}(\mathcal{O})\times{\mathcal{U}}(\mathcal{O}))\rtimes C_{2}$. The largest imprimitive subgroup which is order-admissible, is $(C_{2}\times C_{2})\rtimes C_{2}$, which indeed appears as a subgroup of a spanning group.
@@ 2. Step 2:
Step $1$ was realised by Eisele–Kiefer–Van Gelder in [EKVG, Table 2]. The sets $\mathcal{S}(\operatorname{M}_{2}(\mathcal{O}))$ can also be read from Table˜B in Appendix˜A. We now deal with Step $2$.
@@ 7.4.3. Possible subgroups which are Frobenius complements
The isomorphism types in Proposition˜7.33 arise thanks to the following description of finite subgroups of a quaternion algebra, which follows from [Voight, Prop. 32.4.1, Lemma 32.6.1 & Prop. 32.7.1].
Lemma 7.35 ([Voight]).
@@
Another important ingredient for Proposition˜7.33 are the following restrictions on the possible order of elements in an exceptional matrix algebra obtained in [EKVG] (see also [EricAngel1, Proposition 12.1.1] for the version below).
Theorem 7.36 ([EKVG]).
@@
Suppose that $H$ is not cyclic. Suppose $H$ is imprimitive, and hence of the form given by Lemma˜7.28, i.e. $H\cong(\Gamma\times\Gamma)\rtimes C_{2}$ or $H\cong\Gamma_{1}\times\Gamma_{2}$, for finite $\Gamma,\Gamma_{1},\Gamma_{2}\leqslant{\mathcal{U}}(\mathcal{O})$. Since $H$ spans a division algebra by assumption, it must be a Frobenius complement. Recall that a Frobenius complement does not have a subgroup of the form $C_{p}\times C_{p}$ for $p$ prime, see [PassPerm, Theorem 18.1]. Thus if $H$ is a Frobenius complement, it must be of the form $\Gamma_{1}\times\Gamma_{2}$ with the orders $|\Gamma_{i}|$ for $i=1,2$ relatively prime and $\Gamma_{i}\leqslant{\mathcal{U}}(\mathcal{O})$. In that case $\Gamma_{1}\times\Gamma_{2}$ is indeed a Frobenius complement by [EricAngel1, Exercise 11.4.3]. Now recall (7.10) which asserts in terms of $A$ when $\mathcal{S}(\mathcal{O})$ is non-empty. For these values the unit groups are recorded in (7.11) and (7.12). From these it follows immediately that the only way for $H$ to be non-cyclic and of the form $\Gamma_{1}\times\Gamma_{2}$, is that some $\Gamma_{i}$ is trivial, which finishes the proof of statement (1).
@@
In [EricAngel1, Example 3.5.7], the strong Shoda pairs $(\Gamma,K)$ are computed for a generalised quaternion group $Q_{4n}$. Recall that in Theorem˜3.7, we show for which $e=e(Q_{4n},\Gamma,K)$, the associated representation $\psi\colon Q_{4n}\to{\mathbb{Q}}Q_{4n}e$ is faithful, namely when
@@
But from [EricAngel1, Example 3.5.7] it is then immediate that the only SSP for $Q_{4n}=\langle x,y\mid x^{2n}=1,x^{n}=y^{2},yx=x^{-1}y\rangle$ which yields a faithful representation, is $(\Gamma,K_{2n})$, with $\Gamma=\langle x\rangle$, and $K_{2n}=\langle x^{2n}\rangle=\{1\}$. Moreover in the authors in loc. cit. [^20] Note that there is a small typo, $A_{d}$ should read $\left(\frac{\alpha_{d}^{2}-4,-1}{{\mathbb{Q}}(\alpha_{d})}\right)$ rather than $\left(\frac{\alpha_{d}^{2}-4,-1}{{\mathbb{Q}}}\right)$, as seen in eq. (3.5.1) right below it., show that the associated component ${\mathbb{Q}}Q_{4n}e$ is given by $\left(\frac{\alpha_{2n}^{2}-4,-1}{{\mathbb{Q}}(\alpha_{2n})}\right)$, where $\alpha_{2n}=\zeta_{2n}+\zeta_{2n}^{-1}$.
@@
For the second part of (3), as well as (4), we use the tables of finite spanning subgroups of exceptional matrix algebras provided by [EricAngel1, Corollary 12.12, Tab. 12.1 and Tab. 12.2]. In particular, in Table˜3 we give, for each exceptional matrix algebra, which groups from the statement occur as a subgroup of some finite spanning subgroup. In Table˜A in Appendix˜A we listed per spanning subgroup the groups from the statement which are isomorphically contained in it, and the content of Table˜3 is then obtained by combining the information in Table˜A and Table˜B.
@@ 7.4.4. Some applications of the classification of finite subgroups
In [Hat] it was proven that there is an isomorphism
@@ Appendix A Tables of groups with a faithful exceptional $2\times 2$ embedding
In this appendix we reproduce [^21] Including the group with SmallGroupID [24, 1], which was accidentally omitted in the original table. [EKVG, Table 2], listing those finite groups $G$ that have a faithful exceptional matrix component (see Definition˜1.1) in the Wedderburn–Artin decomposition of the rational group algebra $\mathbb{Q}G$. In particular, these groups appear as a spanning subgroup of an exceptional $\operatorname{M}_{2}(D)$. For all these groups, we list properties relevant for Proposition˜7.33 in the first table.

### v2

@@ 7.2. Somme illustrative examples
If $\Gamma$ is nilpotent, then it even satisfies the 33rd Zassenhaus conjecture [86, 87]. The only non-nilpotent $\Gamma\in\mathcal{S}(\operatorname{M}_{2}({\mathbb{Z}}[\sqrt{-1}]))$ with $|\Gamma|\neq 48$ or $96$, is the metacyclic group $C_{3}\rtimes C_{8}$ (as can be seen from Table˜B in Appendix˜A). Now note that all strict subgroups of $C_{3}\rtimes C_{8}$ are cyclic. As $C_{3}\rtimes C_{8}$ is metacyclic it satisfies the 11st Zassenhaus conjecture [39, 11]. Moreover by Whitcomb’s theorem it also satisfies the isomorphism problem, thus altogether indeed has the subgroup isomorphism property for all its subgroups, as claimed.
@@
Now the only finite subgroups of $\operatorname{GL}_{2}({\mathbb{Z}}[\sqrt{-1}])$ which are not isomorphic to a subgroup of $\mathrm{U}_{2}({\mathbb{F}}_{3})$ are $D_{12}$ and $Dic_{3}$. Thus by part (1), it remains to show that $He\ncong D_{12}$ or $Dic_{3}$. Now if $He\cong D_{12}$, then this would imply that $V({\mathbb{Z}}[Ge])$, with $Ge\cong\mathrm{U}_{2}({\mathbb{F}}_{3})$, contains a subgroup isomorphic to $D_{6}$. However, the latter is a Frobenius group and thus by [68, Theorem D], $Ge$ would also contain a copy of $D_{6}$, a contradiction. The dicyclic case is treated similarly. ∎
@@ 7.3.1. The setting of components of group rings
By (7.7) we have that $\Phi_{e}(H)$ is a finite subgroup of $V(R[Ge])$, and therefore $|\Phi_{e}(H)|\mid|Ge|$, see for example [18, Corollary 2.7]. To prove (i), we show that $|He|\mid|\Phi_{e}(H)|$. For this, note that (7.8) implies that $He=\pi_{e}(H)=\sigma_{e}(\Phi_{e}(H))$ is an epimorphic image of $\Phi_{e}(H)$ and hence $|He|$ divides $|\Phi_{e}(H)|$, as desired. Since $V(R[Ge])$ has exponent $\exp(Ge)$ by (7.3), it follows moreover that $\exp(\Phi_{e}(H))\mid\exp(Ge)$ (since $\Phi_{e}(H)\leqslant V(R[Ge])$ and the fact that the exponent of a subgroup divides the exponent of the overlying group). But since $He$ is an epimorphic image of $\Phi_{e}(H)$, it follows that $\exp(He)\mid\exp(\Phi_{e}(H))$, which finishes (ii).
@@
Let $G$ be a finite group and $H\leq V({\mathbb{Z}}G)$ such that ${\mathbb{Z}}G={\mathbb{Z}}H$. Then $Ge\cong He$ for every $e\in\operatorname{PCI}({\mathbb{Q}}G)$.
@@ 7.3.3. The setting of simple algebras of reduced degree 2
Let $D$ be an arbitrary finite-dimensional division $\mathbb{Q}$}-algebra. It will be useful to distinguish primitive and imprimitive groups. Recall the following definition (see [5, section 2.1]).
@@
First we consider the case that $\mathcal{Z}(D)={\mathbb{Q}}$. Then $B$ is a simple $\mathcal{Z}(D)$-subalgebra of $\operatorname{M}_{2}(D)$. Therefore, the double centraliser theorem [47, Theorem 2.10.1] implies that $\dim_{{\mathbb{Q}}}(B)$ divides $\dim_{{\mathbb{Q}}}\operatorname{M}_{2}(D)$. The latter in turn is a divisor of $16$ since $\operatorname{M}_{2}(D)$ is exceptional. Now recall that $\dim_{\mathcal{Z}(B)}(B)$ is a square and the reduced degree of $B$ is less than or equal to $2$ by [5, pg 455, Lemma]. Together with the assumptions on $B$, and the fact that
@@ 7.4. Finite subgroups of exceptional simple algebras
Amitsur [2] classified all finite subgroups of a finite-dimensional division algebra. Moreover, for each of these groups, he described an explicit division algebra containing it. From this, one can in particular deduce the possible finite spanning subgroups of a given division algebra. In a highly technical work, Banieqbal [5] described the primitive spanning subgroups of $\operatorname{M}_{2}(D)$. However, in case of an exceptional (simple) matrix algebra, a direct and short proof was provided in [23] for the finite groups spanning such simple algebras. In this section we also describe the non-spanning finite subgroups.
@@ 7.4.1. The imprimitive subgroups
We start by describing the imprimitive groups. For the following statement, we follow the techniques from the proof of [5, Lemma 2.2].
@@
Hence in particular the finite subgroups of $\operatorname{GL}_{2}(D)$ of imprimitive type are easy to describe in terms of finite subgroups of ${\mathcal{U}}(D)$. We will mainly be interested in finite subgroups of $\operatorname{GL}_{2}(\mathcal{O})$. By Lemma˜7.21 and [23], if $\operatorname{M}_{2}(D)$ is exceptional, we have that
@@
Moreover, the above division algebras are left norm euclidean, see [4, Remarks 3.13 & 6.14], and hence have up to conjugation a unique maximal order [12, Proposition 2.9]. The unique maximal order of $\mathbb{Q}$} is $\mathbb{Z}$}, which has unit group $C_{2}=\{\pm 1\}$. For ${\mathbb{Q}}(\sqrt{-d})$ it is its ring of integers, denoted $\mathcal{I}_{d}$ whose unit group is as following, see [4, Section 3]:
@@
This follows from a straightforward case-by-case analysis for each isomorphism type of exceptional components $A$ such that $\operatorname{GL}_{2}(D)={\mathcal{U}}(A)$, combined with the list of spanning subgroups given by [47, Corollary 12.12, Tab. 12.1 and Tab. 12.2] (see also Table˜B in Appendix˜A). By Lemma˜7.28 (see also remark˜7.29), all finite imprimitive subgroups are isomorphically contained in a largest imprimitive subgroup, which is of the form $({\mathcal{U}}(\mathcal{O})\times{\mathcal{U}}(\mathcal{O}))\rtimes C_{2}$. We summarise the maximal imprimitive subgroups which appear as a subgroup of a spanning group in Table˜1 by giving the SmallGroupID of a spanning group containing it. Lemma˜7.28 then allows to deduce all other imprimitive subgroups which are contained in a spanning group. Note that from Table˜1, it follows that only for $\operatorname{M}_{2}(\mathbb{H}_{5})$, there are no order-admissible spanning groups for the largest imprimitive subgroup $({\mathcal{U}}(\mathcal{O})\times{\mathcal{U}}(\mathcal{O}))\rtimes C_{2}$. The largest imprimitive subgroup which is order-admissible, is $(C_{2}\times C_{2})\rtimes C_{2}$, which indeed appears as a subgroup of a spanning group.
@@ 2. Step 2:
Step $1$ was realised by Eisele–Kiefer–Van Gelder in [23, Table 2]. The sets $\mathcal{S}(\operatorname{M}_{2}(\mathcal{O}))$ can also be read from Table˜B in Appendix˜A. We now deal with Step $2$.
@@ 7.4.3. Possible subgroups which are Frobenius complements
The isomorphism types in Proposition˜7.33 arise thanks to the following description of finite subgroups of a quaternion algebra, which follows from [85, Prop. 32.4.1, Lemma 32.6.1 & Prop. 32.7.1].
Lemma 7.35 ([85]).
@@
Another important ingredient for Proposition˜7.33 are the following restrictions on the possible order of elements in an exceptional matrix algebra obtained in [23] (see also [47, Proposition 12.1.1] for the version below).
Theorem 7.36 ([23]).
@@
Suppose that $H$ is not cyclic. Suppose $H$ is imprimitive, and hence of the form given by Lemma˜7.28, i.e. $H\cong(\Gamma\times\Gamma)\rtimes C_{2}$ or $H\cong\Gamma_{1}\times\Gamma_{2}$, for finite $\Gamma,\Gamma_{1},\Gamma_{2}\leqslant{\mathcal{U}}(\mathcal{O})$. Since $H$ spans a division algebra by assumption, it must be a Frobenius complement. Recall that a Frobenius complement does not have a subgroup of the form $C_{p}\times C_{p}$ for $p$ prime, see [69, Theorem 18.1]. Thus if $H$ is a Frobenius complement, it must be of the form $\Gamma_{1}\times\Gamma_{2}$ with the orders $|\Gamma_{i}|$ for $i=1,2$ relatively prime and $\Gamma_{i}\leqslant{\mathcal{U}}(\mathcal{O})$. In that case $\Gamma_{1}\times\Gamma_{2}$ is indeed a Frobenius complement by [47, Exercise 11.4.3]. Now recall (7.10) which asserts in terms of $A$ when $\mathcal{S}(\mathcal{O})$ is non-empty. For these values the unit groups are recorded in (7.11) and (7.12). From these it follows immediately that the only way for $H$ to be non-cyclic and of the form $\Gamma_{1}\times\Gamma_{2}$, is that some $\Gamma_{i}$ is trivial, which finishes the proof of statement (1).
@@
In [47, Example 3.5.7], the strong Shoda pairs $(\Gamma,K)$ are computed for a generalised quaternion group $Q_{4n}$. Recall that in Theorem˜3.7, we show for which $e=e(Q_{4n},\Gamma,K)$, the associated representation $\psi\colon Q_{4n}\to{\mathbb{Q}}Q_{4n}e$ is faithful, namely when
@@
But from [47, Example 3.5.7] it is then immediate that the only SSP for $Q_{4n}=\langle x,y\mid x^{2n}=1,x^{n}=y^{2},yx=x^{-1}y\rangle$ which yields a faithful representation, is $(\Gamma,K_{2n})$, with $\Gamma=\langle x\rangle$, and $K_{2n}=\langle x^{2n}\rangle=\{1\}$. Moreover in the authors in loc. cit. [^20] Note that there is a small typo, $A_{d}$ should read $\left(\frac{\alpha_{d}^{2}-4,-1}{{\mathbb{Q}}(\alpha_{d})}\right)$ rather than $\left(\frac{\alpha_{d}^{2}-4,-1}{{\mathbb{Q}}}\right)$, as seen in eq. (3.5.1) right below it., show that the associated component ${\mathbb{Q}}Q_{4n}e$ is given by $\left(\frac{\alpha_{2n}^{2}-4,-1}{{\mathbb{Q}}(\alpha_{2n})}\right)$, where $\alpha_{2n}=\zeta_{2n}+\zeta_{2n}^{-1}$.
@@
For the second part of (3), as well as (4), we use the tables of finite spanning subgroups of exceptional matrix algebras provided by [47, Corollary 12.12, Tab. 12.1 and Tab. 12.2]. In particular, in Table˜3 we give, for each exceptional matrix algebra, which groups from the statement occur as a subgroup of some finite spanning subgroup. In Table˜A in Appendix˜A we listed per spanning subgroup the groups from the statement which are isomorphically contained in it, and the content of Table˜3 is then obtained by combining the information in Table˜A and Table˜B.
@@ 7.4.4. Some applications of the classification of finite subgroups
In [34] it was proven that there is an isomorphism
@@ Appendix A Tables of groups with a faithful exceptional $2\times 2$ embedding
In this appendix we reproduce [^21] Including the group with SmallGroupID [24, 1], which was accidentally omitted in the original table. [23, Table 2], listing those finite groups $G$ that have a faithful exceptional matrix component (see Definition˜1.1) in the Wedderburn–Artin decomposition of the rational group algebra $\mathbb{Q}G$. In particular, these groups appear as a spanning subgroup of an exceptional $\operatorname{M}_{2}(D)$. For all these groups, we list properties relevant for Proposition˜7.33 in the first table.